# Ramsey Number Counterexample Checking and One Vertex Extension Linearly Related to $s$ and $t$

### v3

@@ Abstract
The Ramsey number $R(s,t)$ is the smallest integer $n$ such that all graphs of size $n$ contain a clique of size $s$ or an independent set of size $t$. $\mathcal{R}(s,t,n)$ is the set of all counterexample graphs without this property for a given $n$. We prove that if a graph $G_{n+1}$ of size $n+1$ has $\max\{s,t\}+1$ subgraphs in $\mathcal{R}(s,t,n)$, then $G_{n+1}$ is in $\mathcal{R}(s,t,n+1)$. Based on this, we introduce algorithms for one-vertex extension and counterexample checking with runtime linearly bound by $s$ and $t$. We prove the utility of these algorithms by verifying $\mathcal{R}(4,6,36)$ and $\mathcal{R}(5,5,43)$ are empty given current sets $\mathcal{R}(4,6,35)$ and $\mathcal{R}(5,5,42)$.
@@ 1 Introduction
Next, we’ll introduce basic observations. Denote $k:=|\mathcal{R}(s,t,n)|$. $k>0\Rightarrow R(s,t)>n$. $k=0$ and $|\mathcal{R}(s,t,n-1)|>0\Rightarrow R(s,t)=n$. The discovery of nonempty counterexample sets has been used extensively to improve lower bounds, with $R(4,6)>35$ and $R(5,5)>42$ being current examples [5, 6]. A full database of known counterexample graphs can be found at [7]. The sets of known counterexamples for $\mathcal{R}(5,5,42)$ and $\mathcal{R}(4,6,35)$ are not provably exhaustive, although McKay and Radziszowski conjecture that $\mathcal{R}(5,5,42)$ is [5].
@@ 1.1 Contribution
We provide and prove our main theorem in Section 2. Using this, we first introduce a new algorithm for counterexample checking in Section 3. This algorithm has a runtime bound linearly to $\max\{s,t\}$. Then, in Section 4, we show a new algorithm for one-vertex extension, or going from $\mathcal{R}(s,t,n)$ to $\mathcal{R}(s,t,n+1)$. Our algorithm has a runtime quadratically bound to $\max\{s,t\}$ instead of exponentially. The algorithm also works independent of choice for $s$, $t$, and $n$. Given an input of a known $\mathcal{R}(s,t,n)$, it generates the set of all $\mathcal{R}(s,t,n+1)$ whose subgraphs of size $n$ are all within the known $\mathcal{R}(s,t,n)$, as opposed to prior algorithms that generate the set of all $\mathcal{R}(s,t,n+1)$ with at least 1 subgraph of size $n$ within the known $\mathcal{R}(s,t,n)$. This limits the utility of the algorithm in searching a broader space but allows for efficient execution. We validate this by verifying $\mathcal{R}(5,5,43)$ and $\mathcal{R}(4,6,37)$ are empty given current sets $\mathcal{R}(5,5,42)$ and $\mathcal{R}(4,6,36)$ on a single PC in under 15 minutes.
@@ 2 Main Theorem
Assume we have a subgraph $G^{\prime}$ of $G_{n+1}$ of size $s$ or $t$. If $G^{\prime}$ excludes any $v_{i}\in S$, then $G^{\prime}$ is a subgraph of the $i$th subgraph in $\mathcal{R}(s,t,n)$. Therefore $G^{\prime}$ cannot be a clique of size $s$ or an independent set of size $t$. If $G^{\prime}$ includes all $v_{i}\in S$, then it is at least of size $\max\{s,t\}+1>\max\{s,t\}$ which is a contradiction. ∎
@@ 3 Counterexample Checking
We can check the belonging of a candidate graph $G_{n+1}$ to counterexamples $\mathcal{R}(s,t,n+1)$ through Algorithm 1.
@@ 4 One-Vertex Extension
We can extend from $\mathcal{R}(s,t,n)$ to $\mathcal{R}(s,t,n+1)$ via Algorithm 2.
@@
This algorithm iterates over at most $2k^{2}n$ candidates. Other extension algorithms look at all ways to connect a new vertex ($2^{n}$) to an existing counterexample, meaning they search $k2^{n}$ candidates. Our value is not always lesser, but allows for the choice of $\min\{2k^{2}n,k2^{n}\}$ candidates when performing extension. Additionally, this algorithm explores all possible graphs with at least 2 subgraphs in $\mathcal{R}(s,t,n)$. Because $\mathcal{R}(s,t,n+1)$ is the set of all graphs with $\max\{s,t\}+1$ subgraphs in $\mathcal{R}(s,t,n)$, we exhaustively find all graphs in $\mathcal{R}(s,t,n+1)$.
Regarding checking counterexamples, we can either use existing algorithms or Algorithm 1. When using Algorithm 1, we have already verified two subgraphs, meaning we only need to run $\max\{s,t\}-1$ additional iterations.
Running Algorithm 2 on a subset of $\mathcal{R}(s,t,n)$, we only generate graphs with at least $\max\{s,t\}+1$ subgraphs in the subset when using Algorithm 1 to check, and at least 2 subgraphs in the subset when checking with other schemes. This is a stricter limitation than other extension algorithms that generate graphs with at least 1 subgraph in the subset.
An additional cost of this algorithm is that, in our implementation, the condition of finding that $G_{n-1}\subset G^{\prime}_{n}$ requires calls to an isomorphism checker.
@@ 5 Results
We wrote a single-threaded python implementation of the high level algorithms provided in Sections 3 and 4. Code was ran on a laptop with a single intel core i7. The code, available at https://github.com/aLehav/ramsey-ove-and-check, is open-sourced and fully documented. Algorithms used and details regarding the code are in Section A. We also attempted to decrement to subsets of $\mathcal{R}(s,t,n-1)$ from given subsets of some $\mathcal{R}(s,t,n)$ to broaden the search space, as done by McKay, Exoo, and Radziszowski [5, 6].
@@ 5.3 Decrementing
We decremented the subset of $\mathcal{R}(4,6,35)$ to a subset of $\mathcal{R}(4,6,34)$ and the subset of $\mathcal{R}(5,5,42)$ to a subset of $\mathcal{R}(5,5,41)$ and tried reconstructing to find new counterexamples. None were found.
@@ 6 Conclusion
We introduce a new theorem and way to look at Ramsey theory counterexamples with regard to smaller counterexamples. We provide algorithms that leverage this, and implement them to verify known results. This new approach may assist in novel methods for one-vertex extension or gluing that yield the complete sets $\mathcal{R}(4,6,36)$, $\mathcal{R}(5,5,43)$, or $\mathcal{R}(3,10,40)$ in the future.
@@ A.1 Algorithms
    In Algorithm 1, we determine if a candidate graph has $\max\{s,t\}+1$ subgraphs of size $n$ that are isomorphic to any one of $k$ counterexamples.
@@
    In Algorithm 2, we find all $k$ counterexamples sharing a subgraph of size $n-1$ with a given candidate graph.
@@
    In Algorithm 2, after finding a counterexample sharing a subgraph of size $n-1$, we connect a vertex to recreate an isomorphism to this counterexample and must check if the new $G_{n+1}$ is a counterexample.
In Algorithm 1, we must determine if a candidate graph has $\max\{s,t\}+1$ subgraphs that are isomorphic to a counterexample. As mentioned, the naive implementation may iterate each of $\max\{s,t\}+1$ arbitrarily chosen subgraphs of $G_{n+1}$ over each of our $k$ counterexamples to check if they are isomorphic. However, by pre-computing the sorted edge degree of each of our $k$ counterexamples, we can save time by only checking a candidate subgraph against the subset of counterexamples with a matching sorted edge degree. We use this hashing to significantly speed up our code, using the count of triangles through each vertex or number of all subgraphs of size 3 through a given vertex lowers the number of counterexamples to check against to 1 or 2 and only adds a runtime component of $O(n^{3})$.
Additionally, if for some $i\in[1,n+1]$, $G_{n+1}-\{v_{i}\}$ is in $\mathcal{R}(s,t,n)$, then any subgraph of $G_{n+1}$ of size $n$ must have a subgraph of size $n-1$ in $\mathcal{R}(s,t,n-1)$. This is because all subgraphs of size $n$ must have an overlapping subgraph of size $n-1$ with $G_{n+1}-\{v_{i}\}$. With this knowledge, we desire mapping $\Psi$ from graphs in $\mathcal{R}(s,t,n-1)$ to sets of neighbors of the $n$th vertex that would produce a graph in $\mathcal{R}(s,t,n)$.
Algorithm 2 could then work by iterating over every graph $G_{n}$ in $\mathcal{R}(s,t,n)$ and every index $i\in[1,n]$. For each set of neighbors in $\Psi(G_{n}-\{v_{i}\})$, we can connect $G_{n}-\{v_{i}\}$ to $v_{n+1}$ in an isomorphically equivalent way. Now, $G_{n+1}$ has two subgraphs verified to be in $\mathcal{R}(s,t,n)$. If for $\max{s,t}-1$ additional indices $j\notin\{i,n+1\}$, $\Psi(G_{n+1}-\{v_{n+1},v_{j}\})$ includes the neighbors of $v_{n+1}$ in $G_{n+1}-\{v_{n+1}\}$, then $G_{n+1}-\{v_{j}\}\in\mathcal{R}(s,t,n)$ and so $G_{n+1}$ is a counterexample.
This approach significantly reduces the number of counterexamples we need to find isomorphisms to with candidate graphs in both checking and extension. With this approach we developed the more detailed Algorithm 3 for one-vertex extension and Algorithm 4 for checking within that context.
@@
As an aside, $\max_{G_{n-1}}\{|\Psi(G_{n-1}|\}\leqslant k$, as in practice this would require all graphs in $\mathcal{R}(s,t,n)$ to share a subgraph of size $n-1$.
@@
We can generate $\Psi$ through Algorithm 5. The reason that this algorithm must find all isomorphisms whereas the above algorithms do not is in cases where we add a vertex from one mapping of a given counterexample but check from another. If we only added one isomorphism to the mapping’s output sets, we may incorrectly say a valid counterexample graph is not a counterexample.
@@ A.2 Runtime
We note triangle counts for a graph of size $n$ runs in $O(n^{3})$. We use VF2++ [9] for isomorphism, making our time to hash and find the isomorphism of a single graph $h:=O(n^{3}+1\cdot VF2++)$. This gives, for an arbitrary hashing scheme, $h=O(s+(C+N-1)\cdot\Phi)$ when the scheme runs in $s$ time and leaves at most $C$ valid counterexamples when we want to find $N$ isomorphisms from the candidate to counterexample and isomorphism generation takes $\Phi$ time for an arbitrary scheme. For dictionary construction, we aim to find the maximum $N$ per graph, and so we’ll denote it $h^{\prime}$ in the following analysis.
If we extend from $\mathcal{R}(s,t,n)$ of size k, we take $O(kh^{\prime})$ to generate $\Psi$, iterate over $2kn\max_{G_{n-1}}\{|\Psi(G_{n-1}|\}=O(k^{2}n)$ candidate graphs, and for each take $O(\max\{s,t\}h)$ time to assign edges and check if a counterexample, for an overall runtime of $O(k^{2}n\max\{s,t\}h+kh^{\prime})$.
@@ Acknowledgments
Thanks to Afia Anjum for her endless patience in listening to me ramble about these theorems.

### v4

# Ramsey Number Counterexample Checking and One Vertex Extension Linearly Bound by $s$ and $t$

@@ Abstract
The Ramsey number $R(s,t)$ is the smallest integer $n$ such that all graphs of size $n$ contain a clique of size $s$ or an independent set of size $t$. $\mathcal{R}(s,t,n)$ is the set of all counterexample graphs without this property for a given $n$. I prove that if a graph $G_{n+1}$ of size $n+1$ has $\max\{s,t\}+1$ subgraphs in $\mathcal{R}(s,t,n)$, then $G_{n+1}$ is in $\mathcal{R}(s,t,n+1)$. Based on this, I introduce algorithms for one-vertex extension and counterexample checking with runtime linearly bound by $s$ and $t$. The utility of these algorithms is shown by verifying $\mathcal{R}(4,6,36)$ and $\mathcal{R}(5,5,43)$ are empty given current sets $\mathcal{R}(4,6,35)$ and $\mathcal{R}(5,5,42)$.
@@ 1 Introduction
Some basic observations are as follows: Denote $k:=|\mathcal{R}(s,t,n)|$. $k>0\Rightarrow R(s,t)>n$. $k=0$ and $|\mathcal{R}(s,t,n-1)|>0\Rightarrow R(s,t)=n$. The discovery of nonempty counterexample sets has been used extensively to improve lower bounds, with $R(4,6)>35$ and $R(5,5)>42$ being current examples [5, 6]. A full database of known counterexample graphs can be found at [7]. The sets of known counterexamples for $\mathcal{R}(5,5,42)$ and $\mathcal{R}(4,6,35)$ are not provably exhaustive, although McKay and Radziszowski conjecture that $\mathcal{R}(5,5,42)$ is [5].
@@ 1.1 Contribution
I provide and prove the main theorem in Section 2. Using this, I first introduce a new algorithm for counterexample checking in Section 3. This algorithm has a runtime bound linearly by $\max\{s,t\}$. Then, in Section 4, I show a new algorithm for one-vertex extension, or going from $\mathcal{R}(s,t,n)$ to $\mathcal{R}(s,t,n+1)$. The algorithm has a runtime linearly bound by $\max\{s,t\}$ instead of exponentially. The algorithm also works independent of choice of $s$, $t$, and $n$. Given an input of a known $\mathcal{R}(s,t,n)$, it generates the set of all $\mathcal{R}(s,t,n+1)$ with at least 2 subgraphs within the known $\mathcal{R}(s,t,n)$, as opposed to prior algorithms that generate the set of all $\mathcal{R}(s,t,n+1)$ with at least 1 subgraph within the known $\mathcal{R}(s,t,n)$. This limits the utility of the algorithm in searching a broader space but allows for efficient execution. I validate this by verifying $\mathcal{R}(5,5,43)$ and $\mathcal{R}(4,6,37)$ are empty given current sets $\mathcal{R}(5,5,42)$ and $\mathcal{R}(4,6,36)$ on a single PC in under 15 minutes.
@@ 2 Main Theorem
Assume there exists a subgraph $G^{\prime}$ of $G_{n+1}$ of size $s$ or $t$. If $G^{\prime}$ excludes any $v_{i}\in S$, then $G^{\prime}$ is a subgraph of the $i$th subgraph in $\mathcal{R}(s,t,n)$. Therefore $G^{\prime}$ cannot be a clique of size $s$ or an independent set of size $t$. If $G^{\prime}$ includes all $v_{i}\in S$, then it is at least of size $\max\{s,t\}+1>\max\{s,t\}$ which is a contradiction. ∎
@@ 3 Counterexample Checking
Algorithm 1 checks the belonging of a candidate graph $G_{n+1}$ to counterexamples $\mathcal{R}(s,t,n+1)$.
@@ 4 One-Vertex Extension
Algorithm 2 extends from $\mathcal{R}(s,t,n)$ to $\mathcal{R}(s,t,n+1)$.
@@
This algorithm iterates over at most $2k^{2}n$ candidates. Other extension algorithms look at all ways to connect a new vertex ($2^{n}$) to an existing counterexample, meaning they search $k2^{n}$ candidates. Both extension algorithms in conjunction allow for the improved choice of $\min\{2k^{2}n,k2^{n}\}$ candidates. Additionally, this algorithm explores all possible graphs with at least 2 subgraphs in $\mathcal{R}(s,t,n)$. Because $\mathcal{R}(s,t,n+1)$ is the set of all graphs with $\max\{s,t\}+1$ subgraphs in $\mathcal{R}(s,t,n)$, this exhaustively find all graphs in $\mathcal{R}(s,t,n+1)$.
Regarding checking counterexamples, existing algorithms or Algorithm 1 can be used. When using Algorithm 1, only $\max\{s,t\}-1$ additional iterations are needed as two subgraphs have already been verified.
Running Algorithm 2 on a subset of $\mathcal{R}(s,t,n)$, only graphs with at least $\max\{s,t\}+1$ subgraphs in the subset are generated when using Algorithm 1 to check, and at least 2 subgraphs in the subset when checking with other schemes. This is a stricter limitation than other extension algorithms that generate graphs with at least 1 subgraph in the subset.
An additional cost of this algorithm is that, in my implementation, the condition of finding that $G_{n-1}\subset G^{\prime}_{n}$ requires calls to an isomorphism checker.
@@ 5 Results
I wrote a single-threaded python implementation of the high level algorithms provided in Sections 3 and 4. Code was ran on a laptop with a single intel core i7. The code, available at https://github.com/aLehav/ramsey-ove-and-check, is open-sourced and fully documented. Algorithms used and details regarding the code are in Section A. I also decremented to subsets of $\mathcal{R}(s,t,n-1)$ from given subsets of some $\mathcal{R}(s,t,n)$ to broaden the search space, as done by McKay, Exoo, and Radziszowski [5, 6].
@@ 5.3 Decrementing
I decremented the subset of $\mathcal{R}(4,6,35)$ to a subset of $\mathcal{R}(4,6,34)$ and the subset of $\mathcal{R}(5,5,42)$ to a subset of $\mathcal{R}(5,5,41)$ and tried reconstructing to find new counterexamples. None were found.
@@ 6 Conclusion
I introduce a new theorem and way to look at Ramsey theory counterexamples with regard to smaller counterexamples. I provide algorithms that leverage this, and implement them to verify known results. This new approach may assist in novel methods for one-vertex extension or gluing that yield the complete sets $\mathcal{R}(4,6,36)$, $\mathcal{R}(5,5,43)$, or $\mathcal{R}(3,10,40)$ in the future.
@@ A.1 Algorithms
    Algorithm 1 determines if a candidate graph has $\max\{s,t\}+1$ subgraphs of size $n$ that are isomorphic to any one of $k$ counterexamples.
@@
    Algorithm 2 finds all $k$ counterexamples sharing a subgraph of size $n-1$ with a given candidate graph.
@@
    In Algorithm 2, after finding a counterexample sharing a subgraph of size $n-1$, the algorithm connects a vertex to recreate an isomorphism to this counterexample and checks if the new $G_{n+1}$ is a counterexample.
Algorithm 1 must determine if a candidate graph has $\max\{s,t\}+1$ subgraphs that are isomorphic to a counterexample. As mentioned, the naive implementation may iterate each of $\max\{s,t\}+1$ arbitrarily chosen subgraphs of $G_{n+1}$ over each of the $k$ counterexamples to check if they are isomorphic. However, by pre-computing the sorted edge degree of each of the $k$ counterexamples, time is saved by only checking a candidate subgraph against the subset of counterexamples with a matching sorted edge degree. This hashing significantly speeds up the code. Using the count of triangles through each vertex or number of all subgraphs of size 3 through a given vertex lowers the number of counterexamples to check against to 1 or 2 and only adds a runtime component of $O(n^{3})$ per candidate.
Additionally, if for some $i\in[1,n+1]$, $G_{n+1}-\{v_{i}\}$ is in $\mathcal{R}(s,t,n)$, then any subgraph of $G_{n+1}$ of size $n$ must have a subgraph of size $n-1$ in $\mathcal{R}(s,t,n-1)$. This is because all subgraphs of size $n$ must have an overlapping subgraph of size $n-1$ with $G_{n+1}-\{v_{i}\}$. With this knowledge, we compose a mapping $\Psi$ from graphs in $\mathcal{R}(s,t,n-1)$ to sets of neighbors of the $n$th vertex that would produce a graph in $\mathcal{R}(s,t,n)$. This is done via Algorithm 5, described later.
Algorithm 2 could then work by iterating over every graph $G_{n}$ in $\mathcal{R}(s,t,n)$ and every index $i\in[1,n]$. For each set of neighbors in $\Psi(G_{n}-\{v_{i}\})$, $G_{n}-\{v_{i}\}$ can be connected to $v_{n+1}$ in an isomorphically equivalent way. Now, $G_{n+1}$ has two subgraphs verified to be in $\mathcal{R}(s,t,n)$. If for $\max{s,t}-1$ additional indices $j\notin\{i,n+1\}$, $\Psi(G_{n+1}-\{v_{n+1},v_{j}\})$ includes the neighbors of $v_{n+1}$ in $G_{n+1}-\{v_{n+1}\}$, then $G_{n+1}-\{v_{j}\}\in\mathcal{R}(s,t,n)$ and so $G_{n+1}$ is a counterexample.
This approach significantly reduces the number of counterexamples needed to find isomorphisms to with candidate graphs in both checking and extension. With this approach I developed the more detailed Algorithm 3 for one-vertex extension and Algorithm 4 for checking within that context.
@@
As an aside, $\max_{G_{n-1}}\{|\Psi(G_{n-1}|\}\leqslant k$. It is usually far less than $k$, as in practice this requires all graphs in $\mathcal{R}(s,t,n)$ to share a subgraph of size $n-1$.
@@
Algorithm 5 generates $\Psi$. The reason that this algorithm must find all isomorphisms whereas the above algorithms do not is in cases where a vertex is connected via one mapping of a given counterexample but checked from another. If only one isomorphism is added to the mapping’s output sets, a valid counterexample graph may incorrectly be labeled as a false negative.
@@ A.2 Runtime
Triangle counts for a graph of size $n$ runs in $O(n^{3})$. I use VF2++ [9] for isomorphism, making my time to hash and find the isomorphism of a single graph $h:=O(n^{3}+1\cdot VF2++)$. This gives, for an arbitrary hashing scheme, $h=O(s+(C+N-1)\cdot\Phi)$ when the scheme runs in $s$ time and leaves at most $C$ valid counterexamples when finding $N$ isomorphisms from the candidate to counterexample and isomorphism generation takes $\Phi$ time for an arbitrary scheme. For dictionary construction, I aim to find the maximum $N$ per graph, and so denote it $h^{\prime}$ in the following analysis.
If I extend from $\mathcal{R}(s,t,n)$ of size k, it takes $O(kh^{\prime})$ to generate $\Psi$, iterate over $2kn\max_{G_{n-1}}\{|\Psi(G_{n-1}|\}=O(k^{2}n)$ candidate graphs, and for each take $O(\max\{s,t\}h)$ time to assign edges and check if a counterexample, for an overall runtime of $O(k^{2}n\max\{s,t\}h+kh^{\prime})$.
@@ Acknowledgments
Thanks to Afia Anjum for her endless patience in listening to me talk about these algorithms.